# Flexible Physical Camouflage Generation Based on a Differential Approach

### v1

@@ Abstract
This study introduces a novel approach to neural rendering, specifically tailored for adversarial camouflage, within an extensive 3D rendering framework. Our method, named FPA, goes beyond traditional techniques by faithfully simulating lighting conditions and material variations, ensuring a nuanced and realistic representation of textures on a 3D target. To achieve this, we employ a generative approach that learns adversarial patterns from a diffusion model. This involves incorporating a specially designed adversarial loss and covert constraint loss to guarantee the adversarial and covert nature of the camouflage in the physical world. Furthermore, we showcase the effectiveness of the proposed camouflage in sticker mode, demonstrating its ability to cover the target without compromising adversarial information. Through empirical and physical experiments, FPA exhibits strong performance in terms of attack success rate and transferability. Additionally, the designed sticker-mode camouflage, coupled with a concealment constraint, adapts to the environment, yielding diverse styles of texture. Our findings highlight the versatility and efficacy of the FPA approach in adversarial camouflage applications.
1 Introduction
@@
Previous research has validated the vulnerability of the neural network. Especially in the work of computer vision[27, 61], natural language processing[28], and deep reinforcement learning[25, 26]. While most of them conduct the attack in the digital space, tasks e.g., facial recognition[60, 55], image classification[36, 35], etc., by crafting specially designed perturbation, adversarial patches, etc. Though the patch can be printed out to deploy the physical attack, environmental factors such as view angle, light, hue, etc., deeply affect the attack. Therefore, it is more difficult to conduct an adversarial attack on the model in the physical world as the complex physical constraints (camera pose, etc.).
@@
Recently, the camouflage attack has sprung out by optimizing the adversarial pattern which can be attached to the entire body of the target (e.g., individuals, vehicles, etc.). Different from the patched attack, the camouflage attack needs to take care of an all-round attack and non-rigid structures on the target surface. Therefore, the two most critical aspects of a camouflage attack involve the method of texture generation and the rendering technology applied to the target surface. According to existing works, there are primarily two methods for texture generation: One approach is based on optimizing texture images, achieving full vehicle camouflage through repeated texture overlays[57, 53]. However, such camouflage struggles to effectively deceive target detectors when facing different distances and viewing angles. The other method involves optimizing the entire area of the target surface that requires texture coverage as a whole[48, 47]. Each part of the texture corresponds to a specific region on the target surface, providing better robustness against changes in viewpoint and occlusion. However, this approach increases the size of the texture, resulting in longer rendering times and making optimization more challenging. Another challenge in camouflage attacks is the rendering of textures onto the target surface. Because 3D targets are non-planar, simulating the coverage status in real-world environments becomes more difficult. To address this issue, some works use neural rendering to map textures[47], but they cannot simulate physical factors’ interference. Another line of work, such as DTA[41] and ACTIVE[42], designs differentiable neural renderers that can learn various scene attributes, enabling the generation of adversarial camouflage in a black-box manner. However, it’s worth noting that neural renderers demand extensive data for learning rendering, requiring substantial labeled data. Furthermore, the training data cannot be changed dynamically during the training process, leading to a lack of inherent generalization in the generated texture. In all the aforementioned works, the camouflage covering the target surface cannot be unfolded in UV maps. Consequently, additional processing of textures is required before physical deployment, which increases the risk of adversarial information loss, extends actual deployment time, and lacks flexibility.
@@
In response to the aforementioned challenges and in pursuit of an approach that allows for an adversarial in a sticker manner, enabling the selection of specific areas to cover with adversarial camouflage, we introduce a novel framework, termed Flexible Physical-camouflage Attack(FPA). Diverging from traditional neural rendering approaches, we fashioned a renderer for adversarial camouflage by leveraging a differentiable mesh renderer framework. Within this framework, we harnessed modular differentiable rendering functions to simulate authentic lighting conditions and material variations, ensuring a nuanced representation of the 3D target’s textures. This approach took into account variables such as lighting and material properties, contributing to a more realistic rendering outcome. Then, we learn the adversarial pattern from the diffusion model in a generative way and combine the designed adversarial loss and covert constraint loss to guarantee the adversarial camouflage’s adversarial and covert nature in the physical world. Meanwhile, the camouflage in sticker mode can easily cover the target in the physical world without losing too much adversarial information. The contribution of this work can be summarized as follows:
@@
  • We established a rendering system rooted in a robust 3D framework, sidestepping the intricacies associated with training neural renderers and achieving heightened rendering efficiency. Meanwhile, in combination with the Blender tool, it can generate adversarial camouflage in the form of stickers, minimizing the loss of adversarial information for physical deployment and enhancing the flexibility for application in the physical world.
@@
  • We learn adversarial patterns from diffusion models, and the learned adversarial patterns are more realistic. This approach enables us to expedite the optimization process when generating higher-resolution adversarial camouflage while ensuring clearer camouflage textures.
@@ 2 Related Works
2.1 Physical Adversarial Patch
@@
2.2 Physical Adversarial Camouflage
@@
3 Flexible Physical Camouflage Generation
The difficulty of current adversarial camouflage generation is to map the 2D texture to the 3D object in a differential way. In the meantime, it is difficult to transfer the generated adversarial texture into the physical world in an alignment manner. To solve these problems, this paper proposed a new framework as previously mentioned FPA.
The structure of the work can be depicted as follows, the 3D object rendering tool we applied is a differentiable mesh renderer in which we only need to optimize the 2D texture to obtain the camouflage. Different from previous work, the structure we applied generated with the environment which can be changed dynamically with the background color style.
To achieve differential optimization of textures, we leverage the differentiable mesh renderer111www.pytorch3d.org to render the target object in a specified environment. The rendered image is then fed into object detection models such as YOLOv5[22], YOLOv3[38], Faster R-CNN[39], etc. In contrast to previous approaches where the texture mask is applied directly to the 3D object in a specific view, we manually define the area to be optimized using Blender. This involves selecting faces in the 3D object files that require optimization, followed by a baking procedure to map the camouflage area to a new UV map. Simultaneously, a new texture is generated post-baking, serving as the texture mask for subsequent optimization steps.
To maintain differentiability throughout the process, the differentiable mesh renderer is employed to render the texture onto the 3D object file. The renderer and associated components support batched inputs, enabling the rendering of a batch of output images in a single forward pass. The background-free nature of the rendered image from PyTorch3D poses a challenge in ensuring generalization. To address this, a fusion method is applied to seamlessly integrate the rendered object into diverse landscapes, enhancing generalization. Throughout the entire process, excluding camouflage face selection, the optimization of the adversarial texture follows an end-to-end approach, facilitated by the differentiable rendering procedure in the renderer. This framework allows for versatile texture generation methods, including gradient-based techniques (e.g., FGSM[16], PGD[34]), generative-based models (e.g., GAN[15]), and diffusion-based models, ensuring flexibility and consistency in the overall approach.
Figure 1: The workflow of the proposed framework, the red line denotes the gradient flow.
3.1 Camouflage Faces Selection
As we described before, the neural renderer usually applies the repeated manner in the camouflage pattern covering. And it is difficult to control the place or points that the camouflage attacked on. In this work, before the optimization, the Blender is applied to select the points and faces we want to cover, then the UV mapping is applied to obtain the 2D texture which is the sticker style that can stick to the target in the physical world. This procedure is achieved through texture baking which creates a new texture based on the faces selected. The texture we selected is illustrated in Figure 2.
(a) UV map of selected texture.
(b) Mask image of the texture.
Figure 2: The selected camouflage area on the UV map.
3.2 Camouflage Texture Generation
Different from previous work, as to texture can be mapped to the target 3D object directly, therefore, in this paper, the adversarial texture of the target model can be optimized directly in the 2D space, i.e., the UV map of the texture. To ensure the texture can obtain the adversarial information and can deceive the target model, which means we can transfer the gradient among the synthetic target image, the 3D object, and the UV map, the differentiable mesh renderer is applied, and as described in section 3, the gradient information is transferring.
3.2.1 Generation with Diffusion Model
@@
As there is a differential optimization flow for the texture generation, the diffusion-based mode can be applied to generate the adversarial texture and suppose the input is noise data $\epsilon$ which is a Gaussian distribution $N(0,I)$ generally, then the generator with U-Net as the backbone to generate the image $T_{k}$ at time $k$, then each diffusion step can be described as follows: The generator takes the image $T_{k}$ and adds Gaussian noise $\epsilon_{k}$ to it, where $\epsilon_{k}$ is a Gaussian distribution $N(0,\sigma_{k}^{2}I)$, and $\sigma_{k}$ is the standard deviation of the Gaussian noise at time $k$, which is shown as follows:
$T_{k-1}=G_{\theta}(T_{k}+\epsilon_{k})$ (1)
@@
where $G_{\theta}$ is the generator with U-Net as the backbone, $\theta$ represents the parameters of the generator, and $k$ represents the time step. The generator then applies a series of convolutional operations to the noisy image to obtain a new image $T_{k-1}$.
3.2.2 Generation with One-Step Backward
Apart from the diffusion model to generate the adversarial texture, it can also be used one step backward to optimize the adversarial texture same as the FGSM or PGD, which injects isotropic noise into manipulated samples and using the gradients of the neural network to create adversarial examples. The optimization involves minimizing the difference between the adversarial texture and the original texture, subject to a constraint on the perturbation magnitude or perceptual sensitivity of the adversarial noise. Similarly, the texture $T$ is the trainable parameter, then the gradient is added to $T$ to create the perturbation. Different from the FGSM which needs to ensure the generated perturbation is small enough to be imperceptible to the human eye, the constraints of the texture include natural, similar to the background, and smooth. Therefore, the gradient is added directly in an efficient way, and this procedure can be described as follows:
$\eta={v}\nabla_{T}L(w,y,T_{a})$ (2)
Then the texture is updated with $T_{a}=T_{a}+\eta$. After the one-step updating, the new texture $T_{a}$ will be deferentially covered towards the target with the differentiable mesh renderer, and then the rendered image is fed into the detection model to validate whether it is adversarial or not.
3.3 Differential Rendering
@@
The texture maps optimized through the diffusion model need to be rendered onto the corresponding regions of the 3D target for further optimization. One difficulty here is that almost all current 3D neural render methods can only process one texture, that is to say, there should only be one texture for a 3D model, otherwise, it can not be correctly rendered. To cope with this problem, the texture blending technique is applied to map different textures in a single 3D object. The texture blending technique involves blending multiple textures based on the blending factor. The blending factor is the texture index associated with each vertex. Then the linear interpolation is applied to fuse the texture, which can be expressed as $T_{blended}=\sum_{i}^{K}\lambda_{i}\cdot T_{i}$, where $T_{belended}$ is the resulting blended texture, and $T_{i}$ is the original texture, and $\lambda_{i}$ is the blending factor which is 1 in the general case.
Therefore, we utilize a texture blending approach to replace the corresponding regions in the original texture with baked textures. After the differential rendering with differentiable mesh renderer $R$, the fusion step is applied to join the background and the rendered object in the same picture to create the new scenario, and this procedure can be described as follows:
$x_{adv}=R(T_{belended},M,P,E)\cdot m_{bk}+x\cdot(1-m_{bk}),$ (3)
where $M$ represents the model file of the target being attacked, $P$ represents the positional information between the target and the camera, $E$ represents environmental information (including brightness, lighting intensity, reflection parameters, etc.), and $m_{bk}$ represents the mask image segmented out from the target (with the segmented area set to 1 and the background area set to 0). The algorithm is described in Algorithm 1. Then, the rendered batch images will be treated as training data to obtain the adversarial texture.
@@
3:  Load the background images $x$, the mask images $m_{bk}$,
4:  Obtain the texture tensor $T$ with “maps padded”
5:  Replace the stacked baked textures in $T$ with adversarial textures $T_{k}$ generated by the diffusion model.
6:  Create a Phong renderer by composing a rasterizer and a shader. The textured Phong shader will interpolate the texture UV coordinates for each vertex, sample from a texture image, and apply the Phong lighting model, set parameters $E$
7:  Set the range of the elevation, azimuth, and distance between the object and the camera
8:  for i in batch size do
9:     Sample the elevation, azimuth, and distance from the defined range, denoted as parameters $P$
10:     Render the image with the defined elevation, azimuth, and distance.
11:     Obtain the render images with Equation 3
12:  end for
3.4 Loss constraints
3.4.1 Adversarial Loss
After the object detector identifies the target, it outputs predicted confidence scores and the bounding box positions. Our objective is to evade detection by the detector, thus the designed adversarial loss ${L_{adv}}$ consists of two components: ${L_{iou}}$ and ${L_{obj}}$. Firstly, ${L_{iou}}$ is defined as the intersection over union (IOU) between the bounding boxes output by the object detector and the bounding boxes in the ground truth labels. During the non-maximum suppression process of object detection, bounding boxes with an IOU lower than a threshold are directly filtered out. Therefore, our objective is to reduce the overlap area of bounding boxes, thereby decreasing the IOU value. ${L_{obj}}$ is defined as the confidence score of the bounding boxes containing the target. We identify the bounding boxes with the highest confidence score as the target and optimize towards minimizing it during the training process, thereby suppressing the recognition of the object detector. Therefore, our adversarial loss is defined as follows:
$\begin{cases}L_{iou}=max(IOU(y^{(x)},b_{i}^{(x)})),\\ L_{obj}=max(c^{(x)}),\\ L_{ADV}=min(\alpha L_{iou}+\beta L_{obj})\end{cases},i=1,2,\ldots N.$ (4)

### v3

@@ Abstract
To achieve attacks on object detection in the physical world, many existing studies focus on adversarial camouflage generation techniques, which involve applying adversarial textures onto the surface of the target. However, the neural renderers used in previous works cannot realistically simulate the physical world, and most texture rendering processes are implemented on 2D images, without fully considering the influence of the structural information of 3D models on texture overlay. To address these challenges, we introduce a novel neural rendering approach specifically tailored for adversarial camouflage within a versatile 3D rendering framework. Our method, named Flexible Physical-camouflage Attack (FPA), goes beyond traditional techniques by faithfully simulating lighting conditions, material variations of the camouflage, the angle and distance between the target and the camera, ensuring a nuanced and realistic representation of textures on a 3D target. To achieve this, we employ a generative approach that learns adversarial patterns from a diffusion model. This involves combining specially designed adversarial loss, smoothness loss, non-printable score loss, and concealment constraint loss to ensure the adversarial, printability, and concealable nature of camouflage in the physical world. Furthermore, we showcase the effectiveness of the proposed camouflage in sticker mode, demonstrating its ability to cover the target without compromising adversarial information. Through empirical and physical experiments, FPA exhibits strong performance in terms of attack success rate and transferability. Additionally, the designed sticker-mode camouflage, coupled with a concealment constraint, adapts to the environment, yielding diverse styles of texture. Our findings highlight the versatility and efficacy of the FPA approach in adversarial camouflage applications.
Index Terms:
adversarial camouflage, physical adversarial attacks, security of neural network.
I Introduction
@@
Previous research has confirmed the vulnerability of neural networks, particularly in computer vision [27, 65], natural language processing[28], and deep reinforcement learning[25, 26]. Most of the work involves attacks in the digital space, such as disrupting facial recognition[64, 58], image classification tasks[36, 35], etc., through carefully designed perturbations, adversarial patches. Though the patch can be printed out to deploy the physical attack, environmental factors such as view angle, light, hue, etc., deeply affect the attack. Therefore, it is more difficult to conduct an adversarial attack on the model in the physical world as the complex physical constraints (camera pose, etc.).
@@
Recently, the camouflage attack has sprung out by optimizing the adversarial pattern which can be attached to the entire body of the target (e.g., individuals, vehicles, etc.). Different from the patched attack, the camouflage attack needs to take care of an all-round attack and non-rigid structures on the target surface. Therefore, the two most critical aspects of a camouflage attack involve the method of texture generation and the rendering technology applied to the target surface. According to existing works, there are primarily two methods for texture generation: One approach is based on optimizing texture images, achieving full vehicle camouflage through repeated texture overlays[60, 56, 44]. However, the same texture presents different effects on various surfaces of the vehicle, making it challenging for such camouflage to effectively deceive target detectors when faced with varying distances and perspectives. Another approach is to treat the entire area of the target surface as the optimization objective[51, 50], where each part of the texture corresponds to a specific region on the target’s curved surface. This method provides better robustness in coping with variations in viewpoints and occlusions. However, this approach increases the size of the texture, resulting in longer rendering times and making optimization more challenging. Another challenge in camouflage attacks is the rendering of textures onto the target surface. Because 3D targets are non-planar, simulating the coverage status in real-world environments becomes more difficult. To address this issue, some works use neural rendering to map textures[50, 63]. However, these methods consider the impact of environmental factors on camouflage only after obtaining the 2D rendered image and cannot simulate the interference of physical factors during the rendering process. Another approach, such as DTA[44] and ACTIVE[45], is based on repetitive projection or triplanar mapping, mapping adversarial textures onto targets in 2D images. This is essentially a 2D-to-2D rendering process, which inevitably leads to inaccuracies in texture shapes when dealing with non-planar projections. The aforementioned methods, from texture rendering to the generation of final adversarial samples, are not purely 3D processes. They do not sufficiently utilize the structural information of 3D target models. The final optimization yields adversarial textures covering the entire vehicle. However, due to limitations in the rendering methods, it’s not possible to directly unwrap the adversarial textures on the vehicle’s surface into an image. In actual deployment, the texture can only be applied by cropping the rendered images from fixed angles, which ignores angular transformations on the vehicle’s surface. This approach increases the risk of adversarial information loss and lacks flexibility.
@@
In response to the aforementioned challenges, we are the first to propose a comprehensive, multi-view, differentiable adversarial camouflage generation framework from 2D textures to 3D targets, termed Flexible Physically-camouflage Attack (FPA). Diverging from traditional neural rendering approaches, we fashioned a renderer for adversarial camouflage by leveraging a differentiable mesh renderer framework. Within this framework, we harnessed modular differentiable rendering functions to simulate authentic lighting conditions and material variations, ensuring a nuanced representation of the 3D target’s textures. This approach took into account variables such as lighting and material properties, contributing to a more realistic rendering outcome. Then, we learn the adversarial pattern from the diffusion model in a generative way and combine the designed adversarial loss and covert constraint loss to guarantee the adversarial camouflage’s adversarial and covert nature in the physical world. Meanwhile, the camouflage in sticker mode can easily cover the target in the physical world without losing too much adversarial information. The contribution of this work can be summarized as follows:
@@
  • We established a rendering system rooted in a robust 3D framework, sidestepping the intricacies associated with training neural renderers and achieving heightened rendering efficiency. This framework can randomly set the viewpoint of the 3D target, ambient light intensity, and material properties of the camouflage during the rendering process. It generates adversarial camouflage in the form of UV maps, minimizing the loss of adversarial information during physical deployment and enhancing flexibility in real-world applications.
@@
  • We are the first to propose using diffusion models to generate adversarial textures for 3D targets. We learn adversarial patterns from diffusion models, and the learned adversarial patterns are more realistic. This approach enables us to expedite the optimization process when generating higher-resolution adversarial camouflage while ensuring clearer camouflage textures.
@@ II Related Works
II-A Physical Adversarial Patch
@@
II-B Physical Adversarial Camouflage
@@
III Methodology
In this section, we first briefly describe the process of generating adversarial camouflage for 3D objects, explaining the relationship between adversarial camouflage and the victim model, and outlining the requirements for successfully attacking detectors. Subsequently, we provide a detailed analysis of the various components of the attack framework FPA. The overall framework is as illustrated in Figure. 1.
Figure 1: Workflow of generating robust adversarial textures in FPA framework.
III-A Problem Definition
During the training process of adversarial camouflage, the attacker does not need the network structure and parameters of the target model, but only needs to provide input to the target model and obtain the predicted output. Unlike previous methods for training adversarial camouflage, we only need a small number of background images ${x_{bk}}$, without requiring a large dataset containing cars. Given texture $T$, 3D mesh $M$, and camera parameters $P$ (which include the positional information between the object and the camera), differentiable renderer $R$ can generate rendered images of the 3D object in different positions along with the corresponding mask images, $<x_{ren},x_{mask}>=R(T,M,P)$. In $x_{mask}$, 0 represents the area where the object is located, and 1 represents the background area. By identifying the boundary points of the object area, we can obtain the true label $y$ corresponding to $x_{ren}$. We can merge the rendered image and the background using a fusion function $\Phi(\cdot)$, denoted as follows:
$\Phi(x_{ren},x_{mask},x_{bk})=x_{ren}\cdot x_{mask}+x_{bk}\cdot(1-x_{mask})$ (1)
Given a target detection model $O$, upon inputting the target image, we can obtain the detector’s predicted bounding box information $b=(b_{x},b_{y},b_{w},b_{h})$ (where $b_{x}$ and $b_{y}$ are the coordinates of the center of the predicted bounding box, and $b_{w}$ and $b_{h}$ represent the width and height of the predicted bounding box) and the confidence score $c$, collectively referred to as $<b^{(x)},c^{(x)}>={O}(x)$.
Our goal is to generate a robust adversarial texture that, when applied to the surface of a 3D object model through a differentiable renderer, can produce adversarial samples rendered from different positions. We formulate the generation of adversarial texture as an optimization problem, seeking the optimal adversarial texture that enables the object to evade detection by the target detector, as described below:
$\underset{T_{\text{adv }}}{\arg\min}L\left({O}\left(\Phi\left(R\left(T_{adv},M,P\right),x\right)\right),y\right),$ (2)
where $L(\cdot)$ is the loss function.
III-B Camouflage Texture Generation
Unlike the work of DTA[44] and ACTIVE[45], which repeatedly map optimized adversarial textures onto different surfaces of the target, we obtain the UV map of the 3D target surface texture through UV unwrapping. The UV map is a 2D texture map that can be accurately fitted on the 3D model. Subsequently, we can directly optimize the adversarial texture in the 2D space, i.e., the UV map $T_{UV}$ of the texture. Finally, the optimized adversarial texture is rendered onto the surface of the 3D target through the neural renderer. Therefore, our method achieves targeted optimization for all surfaces, ensuring that each region of the generated adversarial texture corresponds to a specific location on the 3D target surface. This guarantees the best attack effect from every viewpoint. As seen in the work of FCA[50] and RAUCA[63], this approach makes the optimization of adversarial textures challenging, with the texture information being difficult to converge.
@@ III-B Camouflage Texture Generation
Since the emergence of diffusion models, their powerful generative capabilities have led to widespread application in the field of image generation[17, 43, 40]. Therefore, we propose optimizing adversarial textures using a diffusion-based approach. We are the first to introduce the use of diffusion models to generate adversarial textures targeting 3D objects. Assuming the input is noise data $\epsilon$, typically following a Gaussian distribution $N(0,I)$, and the generator, based on U-Net[41] architecture, generates the adversarial texture $T_{k}$ at time $k$, each diffusion step can be described as follows: The generator takes the adversarial texture $T_{k}$ and adds Gaussian noise $\epsilon_{k}$ to it, where $\epsilon_{k}$ is a Gaussian distribution $N(0,\sigma_{k}^{2}I)$, and $\sigma_{k}$ is the standard deviation of the Gaussian noise at time $k$, which is shown as follows:
$T_{k-1}=G_{\theta}(T_{k}+\epsilon_{k}),$ (3)
@@
where $G_{\theta}$ is the generator with U-Net as the backbone, $\theta$ represents the parameters of the generator, and $k$ represents the time step. The generator then applies a series of convolutional operations to the noisy image to obtain a new texture $T_{k-1}$. According to the shape of the UV map $T_{UV}$, obtain the desired adversarial texture $T_{adv}$ from the generated texture $T_{k-1}$, i.e., $T_{adv}=T_{k-1}\cdot T_{UV}$.
Traditional gradient-based generation methods produce adversarial textures that are unclear and highly pixelated, failing to consider the impact of actual physical factors. We propose a method for learning adversarial patterns from diffusion models, which leverage diffusion models to capture various complex spatial relationships and interactions, ensuring the stable generation and clarity of adversarial textures. As shown in Figure. 2, adversarial textures generated using diffusion models are more natural and legible . In the Section. IV-F, we further validate the generative capabilities of the diffusion model.
Figure 2: (a) shows the results generated based on the diffusion model, while (b) shows the results generated based on the gradient method.
III-C Differential Rendering
The difficulty of current adversarial camouflage generation is to map the 2D texture to the 3D object in a differential way. Additionally, it is difficult to transfer the generated adversarial texture into the physical world in an alignment manner. Therefore, we chose to use a neural renderer based on UV mapping to render the adversarial textures onto the surface of the 3D target. In works such as RACUA and FCA, the neural renderers[23] used independent light sources and were unable to render complex environmental features. Existing methods typically simulate environmental characteristics (e.g., EOT[1], Automold[2], EFE[63]) with the aid of external modules after obtaining the rendered images. This process is carried out on 2D images, ignoring the impact of the 3D structure (such as the occlusion of light by different parts of a car’s surface or the reflectivity of the surface). To address these issues, we constructed the neural renderer based on a differentiable mesh111www.pytorch3d.org, which supports the creation of new light sources during rendering and allows modifications to the surface material of the model. Then, in a differentiable manner, we render the adversarial texture onto the 3D target surface. This approach ensures that the adversarial texture is applied directly on the 3D model surface throughout the entire process, from rendering to the output of adversarial examples.
@@ III-C Differential Rendering
Texture blending. The texture maps optimized through the diffusion model need to be rendered onto the corresponding regions of the 3D target for further optimization. One difficulty here is that almost all current 3D neural render methods can only process one texture, that is to say, there should only be one texture for a 3D model, otherwise, it can not be correctly rendered. To cope with this problem, the texture blending technique is applied to map different textures in a single 3D object. The texture blending technique involves blending multiple textures based on the blending factor. The blending factor is the texture index associated with each vertex. The initial texture of the 3D model is $T_{raw}=[T_{1},T_{2},...,T_{n}]$. The blended texture can be represented as $B(T_{raw})=\sum_{i}^{n}\lambda_{i}\cdot T_{i}$, and $\lambda_{i}$ is the blending factor, which is typically set to 1. Therefore, we use a texture blending method to replace the corresponding regions in the original texture with the adversarial texture $T_{adv}$, without affecting the normal display of other areas on the 3D model surface. The final blended texture rendered onto the 3D model surface is described as follows:
$T_{sum}=B(T_{raw},T_{adv})$ (4)
Flexible Renderer. Before generating the rendered images, we need to set the position information $P$ between the target and the camera, as well as the environmental information $E$ (including lighting intensity, surface material reflection parameters, etc.). The position information $P$ includes three parameters: distance, pitch angle, and azimuth angle, ensuring the generation of rendered images from different distances and viewpoints. Unlike previous work, the renderer $R$ built with PyTorch3D allows for the addition of new light sources in the rendering space. We set the parameter $\upsilon_{light}$ to control the light intensity and the parameter $\upsilon_{model}$ to control the lighting mode during the rendering process, including global illumination, directional light, and point light sources. In actual deployment, the camouflage’s different material properties can lead to significant differences in texture appearance. We set the parameter $\upsilon_{material}$ to adjust the material of the camouflage during rendering, primarily controlling its reflectivity to light. Therefore, given the input 3D mesh $M$ and the blended texture $T_{sum}$, we can generate rendered images and corresponding mask images through the renderer, as shown in the following equation:
$<x_{ren},x_{mask}>=R(T_{sum},M,P,E)$ (5)
Subsequently, applying the blending step combines the background and the rendered object in the same image to generate more diverse adversarial examples. Combining this process with the Equation 1, the process can be described as follows:
$x_{adv}=\phi(x_{ren},x_{mask},x_{bk})$ (6)
As shown in Figure. 3, in scenes with direct lighting, compared to commonly used neural renderers, our method considers environmental information during the rendering process and employs UV mapping. This enables us to simulate more realistic environments and achieve more accurate texture projection.
Figure 3: Comparison of rendered images with different textures using the Neural Renderer and the Flexible Renderer.
The entire rendering process is detailed in Algorithm 1.
@@
3:  Load the initial texture $T_{raw}=[T_{1},T_{2},...,T_{n}]$ and adversarial texture $T_{adv}$ generated by the diffusion model
4:  Obtain the blended texture $T_{sum}$ with Equation 4
5:  Create a Phong renderer by composing a rasterizer and a shader
6:  for i in batch size do
7:     Sample the elevation, azimuth, and distance from the defined range, denoted as parameters $P$
8:     Set the values of $\upsilon_{light}$, $\upsilon_{model}$, $\upsilon_{material}$, denoted as parameters $E$
9:     Obtain the render images and mask images with Equation 5
10:  end for
III-D Loss constraints
Adversarial Loss. After the object detector identifies the target, it outputs predicted confidence scores and the bounding box positions. Our objective is to evade detection by the detector, thus the designed adversarial loss ${L_{adv}}$ consists of two components: ${L_{iou}}$ and ${L_{obj}}$. Firstly, ${L_{iou}}$ is defined as the intersection over union (IOU) between the bounding boxes output by the object detector and the bounding boxes in the ground truth labels. During the non-maximum suppression process of object detection, bounding boxes with an IOU lower than a threshold are directly filtered out. Therefore, our objective is to reduce the overlap area of bounding boxes, thereby decreasing the IOU value. ${L_{obj}}$ is defined as the confidence score of the bounding boxes containing the target. We identify the bounding boxes with the highest confidence score as the target and optimize towards minimizing it during the training process, thereby suppressing the recognition of the object detector. Therefore, our adversarial loss is defined as follows:
$\begin{cases}L_{iou}=max(IOU(y^{(x)},b_{i}^{(x)})),\\ L_{obj}=max(c^{(x)}),\\ L_{ADV}=min(\alpha L_{iou}+\beta L_{obj})\end{cases},i=1,2,\ldots N,$ (7)